Direct Materials and Direct Labor Variances

Zoller Company produces a dark chocolate candy bar. Recently, the company adopted the following standards for one bar of the candy:

Direct materials (6.30 oz. @ $0.20) $1.26
Direct labor (0.08 hr. @ $18.00) 1.44
Standard prime cost $2.70

During the first week of operation, the company experienced the following actual results:

  1. Bars produced: 144,000.
  2. Ounces of direct materials purchased: 907,500 ounces at $0.21 per ounce.
  3. There are no beginning or ending inventories of direct materials.
  4. Direct labor: 11,380 hours at $17.30.

Required:

Instructions for parts 1 and 2: If a variance is zero, enter "0" and select "Not applicable" from the drop down box.

1. Compute price and usage variances for direct materials.

Materials Price Variance $
Materials Usage Variance $

2. Compute the rate variance and the efficiency variance for direct labor.

Labor Rate Variance $
Labor Efficiency Variance $

3. Prepare the journal entries associated with direct materials and direct labor. If an amount box does not require an entry, leave it blank or enter "0".

Record purchase of materials
Record usage of materials
Record labor variances
Answer #1

Solution 1:

Direct Material Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AQ * AP = AQ * SP = SQ * SP =
907500 $0.21 $190,575.00 907500 $0.20 $181,500.00 907200 $0.20 $181,440.00
$9,075.00 U $60.00 U
Direct Material Price Variance Direct Material Qty variance
Direct material price variance $9,075.00 U
Direct material quantity variance $60.00 U

Solution 2:

Direct Labor Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
11380 $17.30 $196,874.00 11380 $18.00 $204,840.00 11520 $18.00 $207,360.00
$7,966.00 F $2,520.00 F
Direct Labor rate Variance Direct Labor Efficiency Variance
Direct Labor Rate variance $7,966.00 F
Direct Labor Efficiency variance $2,520.00 F

Solution 3:

Journal Entries - Zoller Company
Event Particulars Debit Credit
1 Raw material inventory Dr $181,500.00
Direct material Price variance Dr $9,075.00
       To Accounts payable $190,575.00
(To record purchase of materials)
2 Work In Process Dr $181,440.00
Direct material quantity variance Dr $60.00
       To Raw material inventory $181,500.00
(To record use of materials)
3 Work In Process Dr $207,360.00
       To Wages Payable $196,874.00
       To Direct labor rate variance $7,966.00
       To Direct labor efficiency variance $2,520.00
(To record labor variances)